Simplify: 3/4/4/5
Possible Answers:
A. 16/15
B. 15/16
C. 5/3
D. 3/5
I am coming up with 3/80, since that is not one of the possible answers then I think I need some assistance.
Thanks

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

(3/4)/(4/5)=3/4*5/4=15/16 ANSWER.

Answer by stanbon(75887)

Simplify: 3/4/4/5
Possible Answers:
A. 16/15
B. 15/16
C. 5/3
D. 3/5
-------------------
Invert the denominator and multiply to get:
(3/4)*(5/4) = 15/16
==========================
Cheers,
Stan H.